MAPPING HYDRATION DYNAMICS AND COUPLED WATER-PROTEIN FLUCTUATIONS AROUND A PROTEIN SURFACE

\underline{LUYUAN ZHANG}, LIJUAN WANG, YA-TING KAO, WEIHONG QIU, YI YANG, OGHAGHARE OKOBIAH, and DONGPING ZHONG, Departments of Physics, Chemistry, and Biochemistry, Programs of Biophysics, Chemical Physics, and Biochemistry, 191 West Woodruff Avenue, The Ohio State University, Columbus, OH 43210.

Elucidation of the molecular mechanism of water-protein interactions is critical to understanding many fundamental aspects of protein science, such as protein folding and misfolding and enzyme catalysis. We recently carried out a global mapping of protein-surface hydration dynamics around a globular α-helical protein apomyoglobin. The intrinsic optical probe tryptophan was employed to scan the protein surface one at a time by site-specific mutagenesis. With femtosecond resolution, we mapped out the dynamics of water-protein interactions with more than 20 mutants and for two states, native and molten globular. A robust bimodal distribution of time scales was observed, representing two types of water motions: local relaxation and protein-coupled fluctuations. The time scales show a strong correlation with the local protein structural rigidity and chemical identity. We also resolved two distinct contributions to the overall Stokes-shifts from the two time scales. These results are significant to understanding the role of hydration water on protein structural stability, dynamics and function.